OpenStudy (anonymous):

tan^-1(ln x^2) what's the derivative

OpenStudy (anonymous):

chain rules ....

OpenStudy (anonymous):

[tan^-1] '(a) * ln'(b) * (x^2)'

OpenStudy (anonymous):

2x/1+ (ln x^2)^2 is what im getting

OpenStudy (anonymous):

hmm tan-1(a) = a'/(1+a^2) , but a = ln(b) a' = b'/b b'/b(1+(ln b)^2), but b=x^2 b' = 2x 2x/x^2(1+(ln x^2)^2) 2/x(1+ (ln x^2)^2)

OpenStudy (anonymous):

you was close

OpenStudy (anonymous):

Okay thanks I always have trouble with the chain rule I need to be more careful

OpenStudy (anonymous):

yes, being careful is key, which is why when they get complicated looking i do like i did ... focus on the parts that are deriving and fill in as i go